Week of 190805

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local: Miro (Chair, DB), Maria (Storage), Jiri (Atlas), Vincent (Security)
  • remote: Xavier (KIT), Luca (LHCb), Marcelo (INFN), Mike (ASGC), Sang-Un (KISTI), Ville (NDGF), Xin (BNL), Dave (FNAL), David (IN2P3), Brian (RAL)

Experiments round table:

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • MC, user jobs and staging.
      • in the coming weeks we will perform a lot of tape recall.
      • remind T2D sites that they should close their solved tickets
    • Issues:

Sites / Services round table:

  • ASGC:
    • 4th August, suffering some routing problem to IN2P3 and FZK (KIT), which cause many transfer time out error
      • We gave the correct routing prefix to CERN
      • It seems to be the IPv6 routing issue on foreign sites
      • Now the connection to FZK is recovered, and now waiting the fix of IN2P3

  • EGI: NC
  • IN2P3: ticket about problem on transfers with ASGC under investigation
  • JINR: High level of CMS data transfer errors to FNAL. Working on it.
  • KIT:
  • NDGF: arcce01.hpc.ku.dk down for couple of days and tape pools too in Copenhagen after thunderstorm last week.
  • NL-T1: NC
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL: FTS upgrade planned for 6th of August.

  • CERN computing services: NTR
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS:
    • The maintenance downtime last Wed had to be extended by a few hours
      • Due to problems with the upgrade of the service back-end (BMC AR Server)
  • Monitoring:
    • NTR
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR


Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2019-08-05 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback